# Evaluating LLMs’ Inherent Multi-hop Reasoning Ability

### v4

@@ 3.2 Evaluation Method
This part is the basis of the whole experiment and all evaluation results. Following reading comprehension [43], evaluation is conducted through lexical matching using two widely used metrics to assess the performance of models. In this section, we employ F1 and EM scores to evaluate the answers to sub-questions, similar to the single-hop QA task.
@@
Table 2 illustrates the differences between our evaluation method and previous evaluation methods of counterfactual QA and factual MHQA datasets. To interpret the behavior of existing LLMs on each hop of the reasoning process required for multi-hop questions, and to determine their reasoning ability to answer simple questions, we followed the experiment setting proposed by [10]. For example, in the 2-hop dataset, each data contains a 2-hop question, 2 sub-questions, 2 intermediate answers, and a final answer. In order to understand whether LLMs can correct answers by following the right reasoning chain, we calculate the proportion of right and incorrect reasoning chains to compare LLMs’ reasoning performance. Each question or sub-question has two results, correct or incorrect, thus an N-hop question with its N sub-questions has $2^{(N+1)}$ different reasoning chains. Due to the space limitation, we measure and collect correctness statistics for the 2-hop question dataset, $q_{sub1}$, $q_{sub2}$, and $q$, and show the percentage of 8 reasoning chains given by LLMs.
The Joint Performance of Sub-QA and Multi-hop QA
The previous MHQA benchmarks are traditionally evaluated on the EM or F1 score on the final answer [43, 5, 6], which is partially correct. The previous MHQA systems and LLMs are treated as a black box and we can not figure out how they find the final answer. Hence, the final answer evaluation shows limitations as it does not consider whether previous MHQA systems could answer sub-questions correctly or not. To understand the impact of sub-qa on MHQA, we introduce a joint performance that combines the evaluation of Sub-QA performance and MHQA performance. The details of computing the joint scores are shown in appendix B.
@@ 4 Experiments
We conduct comprehensive experiments and evaluate different LLMs by using IRE to answer the following questions: 1) Do LLMs show a performance gap between the Wikipedia-based factual MHQA datasets and our synthesized counterfactual MHQA data? 2) When inputting counterfactual questions, how do LLMs perform in terms of their reasoning ability? 3) How do sub-questions affect the performance of LLMs? 4) How do LLMs perform on reasoning chain evaluation? These investigations aim to shed light on the capabilities and limitations of LLMs when dealing with counterfactual MHQA and multi-step reasoning tasks.
4.1 Experiment Settings
We evaluate LLMs on the randomly selected 900 data from Wikipedia-based MHQA datasets (300 HotpotQA [5], 300 2WikiMultihopQA[6], and 300 MuSiQue [20]) and our counterfactual MHQA data (divided into 2-hop, 3-hop, and 4-hop subsets). We employ the proprietary LLMs in our experiments and to enhance reproducibility, we set the temperature to 0 for proprietary models, and all the experiment results are the average scores of three experiment results. For baselines, we adopt GPT-4 [44], GPT-3.5 [45], text-davinci-003, Bing Chat, and GEMINI-pro [46]. To decouple LLMs’ internal memory and reasoning ability and let LLMs retrieve answers from the given passage as much as possible, we design a prompt that requires LLMs to only retrieve answers based on the given context. The prompt of QA is also shown in the Appendix A.
4.2 Experimental Results
@@
The results of the comparison between the selected Wikipedia-based MHQA data and our counterfactual MHQA data can be found in table 3. LLMs show a performance gap between the selected data and ours. Taking GPT-4 as an example, GPT-4 achieves high EM and F1 scores (69.9 and 82.3, respectively), which are even close to well-finetuned small QA models. While for our 2-hop dataset, EM and F1 scores are sharply declined (53.1 and 62.8). For 3-hop and 4-hop datasets, GPT-4 even performs worse. Since our synthesized data is new, unprecedented knowledge, our results objectively reflect the real reasoning performance of LLMs. In light of the results, we can find that LLMs achieve an inflated high performance on the Wikipedia-based MHQA dataset possibly because of the data contamination that leads to utilizing LLMs’ memory ability rather than reasoning ability.
Table 3: Performance gap between Wikipedia-based factual multi-hop QA datasets and our 2-hop, 3-hop, and 4-hop counterfactual MHQA data. The table reveals that LLMs show an obvious performance gap between previous datasets and our data. The performance is measured by EM and F1 scores with a zero-shot setting.
Models GPT-4 GPT-3.5 GEMINI-pro text-davinci-003 Bing Chat Metric EM F1 EM F1 EM F1 EM F1 EM F1 Wiki HotpotQA 69.9±1.5 82.3±1.3 58.6±0.9 69.1±1.1 58.2±1.3 68.4 ±1.3 50.3±0.9 61.4±0.8 68.1 ±0.6 78.3±1.2 2Wiki 59.7±1.4 67.4±2.7 56.3±0.9 67.6±0.8 48.5±1.6 58.5±0.9 42.3±1.4 53.9±1.5 58.9±0.5 69.9±0.5 MuSique 57.3±1.9 65.4±2.9 49.3±0.8 63.2±1.5 41.3±1.5 54.5±0.7 40.2±0.9 51.0±1.3 49.6±1.1 64.1±0.8 IRE 2-hop 53.1±0.4 62.8±1.3 40.6±0.7 56.7±0.5 35.0±0.7 45.3±1.6 32.6±0.9 48.5±0.8 41.9±0.8 53.4±0.9 3-hop 44.5±0.3 56.4±1.7 37.7±0.5 50.9±1.1 29.6±0.5 42.7±0.9 27.8 ±0.9 46.3±0.8 39.6±1.1 49.4±1.2 4-hop 42.3±0.6 53.5±0.9 32.5±1.2 44.6±0.8 26.1±1.1 35.3±1.2 24.8±0.8 44.1±0.9 30.7±0.9 42.2±0.7
@@
Setting 2 hop 3 hop 4 hop EM F1 EM F1 EM F1 GPT-4 w/o Sub-Q 43.8±0.2 65.2±0.3 41.4±0.2 61.6±0.4 38.1±0.5 48.9±0.3 w Sub-Q 53.2±0.5 67.7±0.6 44.5±0.2 64.5±0.3 42.1±0.1 53.1±0.4 GPT-3.5 w/o Sub-Q 34.3±0.2 51.3±0.1 32.7±0.3 48.6±0.3 31.2±0.5 41.7±0.4 w sub-Q 40.4±0.5 56.9±0.4 37.5±0.1 50.2±0.2 33.5±0.2 45.9±0.6 GEMINI-pro w/o Sub-q 25.2±0.5 55.2±0.7 20.8±0.6 38.7±0.4 14.1 ±0.2 31.0±0.4 w sub-Q 34.6±0.5 64.2±0.5 27.3±0.2 42.1±0.2 25.9±0.2 33.8 ±0.3 text-davinci-003 w/o Sub-q 24.1±0.7 48.8±0.3 22.1±0.4 45.6±0.2 20.0 ±0.1 42.1±0.2 w sub-Q 32.3±0.4 52.7±0.4 27.3±0.3 46.4±0.3 24.2±0.5 42.8±0.7 Bing Chat w/o Sub-q 37.2±0.2 62.4±0.5 33.3±0.5 54.2±0.5 29.3±0.4 48.7±0.3 w sub-Q 41.8±0.5 66.8±0.5 40.1±0.2 58.4±0.5 30.4±0.3 49.7±0.3
Table 4: In the ablation study of the MHQA task, we remove the sub-question information from the prompt and only ask LLMs to get the final answer.
@@
Figures in 4 show the performance of LLMs on the different hops of questions. According to the observation of the three figures, we find that with the hop increases, the complexity of multi-hop questions also increases, leading to the LLMs’ performance decrease. Figure 3 shows that LLMs also suffer the error propagation. When incorrectly answering the previous sub-question, the latter one will also be influenced. Consequently, the performance of Sub_Q2 is worse than that of Sub_Q1. Tables 10 and 11 also illustrate the sub-qa performance of LLMs on the 3-hop and 4-hop datasets in appendix C.
2 hop 3 hop 4 hop F1 $RC$ EM $RC$ F1 $RC$ EM $RC$ F1 $RC$ EM $RC$ GPT-4 0.7 1.5 1.1 1.6 2.3 4.2 GPT 3.5 1.7 2.4 2.7 3.2 3.6 5.8 GEMINI-pro 2.1 3.9 4.6 8.7 5.4 9.5 text-davinci-003 2.4 2.9 3.9 5.2 5.5 7.4 Bing Chat 0.9 1.9 4.2 8.4 4.7 8.9
Table 5: LLMs’ joint performance on the whole reasoning chain. The scores are the average scores of three experiment results
@@
In this part, due to space limitations, we calculate the proportion of the reasoning chain on the 2-hop dataset and present the table. We follow the setting of [10] in calculating the percentage of correct or incorrect answers and record the results. Table 6 shows the reasoning chain evaluation results. The green row shows the percentage of examples whose multi-hop questions can be correctly answered from the right reasoning chain. The red rows show the percentage of examples whose multi-hop questions can be correctly answered but through an incorrect reasoning chain. Among these examples, we notice that there is a low percentage of the LLMs successfully getting the correct final answer based on the right reasoning chain. There is also a large proportion of incorrect final answers as shown in rows 2,4,6 and 8. We take the results of GPT-3.5 as an example, the right reasoning chain only accounts for 13.3% although it shows a relatively high QA performance in previous tables. The percentage of incorrect reasoning chain of GPT-3.5 is 17.7% (sum of the three red rows). However, total failure cases account for 69% (sum of rows 2, 4, 6, and 8) which is substantial for the whole dataset. We conclude that LLMs only get a small proportion of the right reasoning chain and their high performance is relatively inflated due to the considerable proportion of incorrect reasoning chain.
@@
The joint F1 $RC$ and joint EM $RC$ scores in table 5 are the whole reasoning chain evaluation results. We find that with the increases in the reasoning chain, the performances of LLMs dropped swiftly. For example, the Bing Chat could get comparable performance with GPT-4 (0.7 joint F1) on answering 2 hop questions, and get a 0.9 joint F1 score. However, in the 3-hop question, the joint F1 $RC$ and joint EM $RC$ scores of Bing Chat are 4.2 and 8.4. In the 4-hop dataset, Bing Chat gets 4.7 joint F1 $RC$ and 8.9 joint EM $RC$ scores respectively. Since the joint performance is a negative log, the larger scores mean the worse performance on the reasoning chain. We can conclude that LLMs’ reasoning ability decreases with the increases in reasoning chain length.
$q_{sub1}$ $q_{sub2}$ $q$ GPT-4 GPT-3.5 GEMINI-pro text Bing Chat c c c 36.3 13.3 15.0 17.3 28.3 c c w 12.3 9.3 9.0 10.7 7.7 c w c 2.0 6.7 5.3 7.7 6.0 c w w 25.3 24.3 14.7 25.0 16.3 w c c 5.7 3.7 5.3 6.7 2.3 w c w 3.7 3.7 5.3 3.7 3.0 w w c 0.3 7.3 13.3 8.7 5. 0 w w w 14.3 31.7 32.3 30.3 31.3
@@
Table 6: Categorical EM statistics (%) of sub-question evaluation for the five LLMs on our 2-hop dataset. Under the first three columns, c stands for correct and w stands for wrong. For example, the third row shows the percentage of questions where models correctly answer both 2-hop questions and the first sub-question but incorrectly answer the second sub-question. We abbreviate text-davinci-003 as text.
4.3 Ablation Study

### v5

@@ 3.2 Evaluation Metrics
This part is the basis of all evaluation results. Following reading comprehension (Rajpurkar et al., 2016), evaluation is conducted through lexical matching using two widely used metrics to assess the performance of models. We employ F1 and EM scores to evaluate the answers to sub-questions, similar to the single-hop QA task.
@@
Table 2 illustrates the differences between our evaluation method and previous evaluation methods of counterfactual QA and factual MHQA datasets. To interpret the behavior of existing LLMs on each hop of the reasoning process required for multi-hop questions, and to determine their reasoning ability to answer simple questions, we followed the experiment setting proposed by (Tang et al., 2021). For example, in the 2-hop dataset, each data contains a 2-hop question, 2 sub-questions, 2 intermediate answers, and a final answer. In order to understand whether LLMs can correct answers by following the right reasoning chain, we calculate the proportion of right and incorrect reasoning chains to compare LLMs’ reasoning performance. Each question or sub-question has two results, correct or incorrect, thus an N-hop question with its N sub-questions has $2^{(N+1)}$ different reasoning chains. Due to the space limitation, we measure and collect correctness statistics for the 2-hop question dataset, $q_{sub1}$, $q_{sub2}$, and $q$, and show the percentage of 8 reasoning chains given by LLMs.
Joint Performance
Previous MHQA benchmarks are traditionally evaluated on the EM or F1 score on the final answer (Rajpurkar et al., 2016; Yang et al., 2018; Ho et al., 2020), which is partially correct. The previous MHQA systems and LLMs are treated as a black box and we can not figure out how they find the final answer. To understand the impact of sub-qa on MHQA, we introduce a joint performance that combines the evaluation of Sub-QA performance and MHQA performance. The details of computing the joint scores are shown in the Appendix B.
[…]
Datasets Wikipedia HotpotQA 2Wiki MuSiQue Metrics EM F1 PM $\dagger$ EM F1 PM $\dagger$ EM F1 PM $\dagger$ Proprietary LLMs GPT-4 69.9±1.5 82.3±1.3 74.8 ±1.2 59.7±1.4 67.4±2.7 64.8±0.9 57.3±1.9 65.4±2.9 63.9±1.4 GPT-3.5 58.6±0.9 69.1±1.1 62.8±0.7 56.3±0.9 67.6±0.8 59.4±0.9 49.3±0.8 63.2±1.5 53.1±0.6 GEMINI-pro 58.2±1.3 68.4±1.3 63.5±0.9 48.5±1.6 58.5±0.9 54.7±1.2 41.3±1.5 54.5±0.7 46.9±1.3 text 50.3±0.9 61.4±0.8 54.9±0.8 42.3±1.4 53.9±1.5 46.7±0.7 40.2±0.9 51.0±1.5 44.6±0.4 Bing Chat 68.1±0.6 78.3±1.2 72.1±1.2 58.9±0.5 69.9±0.5 63.4±0.8 49.6±1.1 64.1±0.8 52.3±0.7 O1-preview 72.2 ±0.6 82.7 ±0.7 76.9 ±1.1 68.7 ±0.9 79.8 ±0.8 72.3 ±1.2 63.9 ±0.9 72.4 ±0.5 67.9±0.8 Open Source LLMs Llama 2-7b 34.5±1.2 41.3±1.1 38.5±0.3 30.6±1.1 34.7±1.1 33.8±0.9 31.7±0.8 35.6±1.2 34.2±0.9 Mistral-7b 30.6±1.5 37.2±1.4 34.9±0.5 27.4±0.6 29.8±0.9 31.4±0.8 25.2±0.7 28.9±0.8 29.2±0.7 Qwen 2-7b 36.2±1.5 43.5±1.3 39.3±0.4 31.7±1.0 35.8±0.8 36.8±0.5 28.2±1.1 31.2±1.2 33.5±0.4
Datasets CofCA 2-hop 3-hop 4-hop Metrics EM F1 PM $\dagger$ EM F1 PM $\dagger$ EM F1 PM $\dagger$ Proprietary LLMs GPT-4 53.1±1.5 62.8±1.3 57.6±1.1 44.5±1.3 56.4±1.7 49.5±1.2 42.3±0.6 53.5±0.9 48.8±1.1 GPT-3.5 40.6±0.7 56.7±0.5 43.7±0.9 37.7±0.5 50.9±1.1 42.1±1.3 32.5±1.2 44.6±0.8 36.2±1.1 GEMINI-pro 35.0±0.7 45.3±1.6 38.2±0.8 29.6±0.5 42.7±0.9 31.9±0.6 26.1±1.1 35.3±1.2 29.8±1.1 text 32.6±0.9 48.5±0.8 37.4±0.9 27.8±0.9 46.3±0.8 33.5±1.2 24.8±0.8 44.1±0.9 27.6±0.7 Bing Chat 41.9±0.8 53.4±0.9 45.4±0.9 39.6±1.1 49.4±1.2 43.7±0.7 30.7±0.9 42.2±0.7 35.6±0.7 O1-preview 59.4 ±0.4 68.5±0.3 63.9±0.8 52.3 ±0.5 66.1±0.4 58.4±0.7 50.7 ±0.4 62.3±0.6 53.2±0.7 Open Source LLMs Llama 2-7b 26.1±0.9 34.3±1.2 30.5±0.4 22.6±1.1 26.7±1.3 25.8±0.8 24.9±1.2 28.7±1.1 30.5±0.9 Mistral-7b 24.7±0.9 29.5±0.7 28.5±0.3 20.8±1.1 25.3±1.1 24.8±0.7 18.6±1.1 22.2±1.3 23.5±0.5 Qwen 2-7b 30.8±1.1 38.2±1.4 34.1±0.4 27.2±1.1 31.7±1.3 31.8±0.4 25.2±0.8 28.6±0.9 29.2±0.5
Table 4: Results on CofCA, the results reveal that LLMs show an obvious performance gap between previous Wikipedia-based factual data and our CofCA.
@@ 4 Experiments
We conduct comprehensive experiments and evaluate different LLMs, using CofCA to answer the following questions: 1) Do LLMs show a performance gap between the Wikipedia-based factual MHQA datasets and our synthesized counterfactual MHQA data? 2) When inputting counterfactual questions, how do LLMs perform in terms of their reasoning ability? 3) How do sub-questions affect the performance of LLMs? 4) How do LLMs perform on reasoning chain evaluation? These investigations aim to shed light on the capabilities and limitations of LLMs when dealing with counterfactual MHQA and multi-step reasoning tasks.
4.1 Experimental Settings
We evaluate LLMs on the CofCA benchmark, including 900 randomly selected data from Wikipedia-based MHQA datasets (300 QA pairs from HotpotQA (Yang et al., 2018), 300 QA pairs from 2WikiMultihopQA(Ho et al., 2020), 300 QA pairs from MuSiQue (Trivedi et al., 2021)), and 900 annotated counterfactual MHQA data (divided into 2-hop, 3-hop, and 4-hop subsets). We employ proprietary LLMs and open-source LLMs in experiments. To enhance reproducibility, we set the temperature to 0 for proprietary models, and all the experiment results are the average scores of three experiment results. We adopt the proprietary LLMs: GPT-4 (Achiam et al., 2023), GPT-3.5 (Ouyang et al., 2022), text-davinci-003, Bing Chat, GEMINI-pro (Team et al., 2023), and Open Source LLMs such as Llama 2-7b, Mistral-7b and Qwen-7b as the baselines. To decouple LLMs’ internal memory and reasoning ability, and let LLMs retrieve answers from the given passage as much as possible, we design a prompt that requires LLMs to only retrieve answers based on the given context. The prompt of QA is shown in the Appendix A.
4.2 Results
@@
The results of the comparison between the selected Wikipedia-based MHQA data and our CofCA can be found in Table 3 and Table 4. LLMs show a performance gap between the selected data and ours. Taking GPT-4 as an example, GPT-4 achieves high EM and F1 scores (69.9 and 82.3, respectively), which are even close to well-finetuned small QA models. While for our 2-hop dataset, EM and F1 scores are sharply declined (53.1 in table 3 and 62.8 in table 4). For 3-hop and 4-hop datasets, GPT-4 even performs worse. Since our synthesized data is new, unprecedented knowledge, our results objectively reflect the real reasoning performance of LLMs. There is also a concern that EM may not be a dependable method due to its constraints in accurately representing real performance, particularly in scenarios where answers involve aliases or abbreviations. For example, the gold answer is "Lionel Messi" and the generated answer is "Messi". The two results should be treated the same but the EM score is not applicable in this scenario. As a result, we introduce GPT-4-turbo as the answer evaluator to measure the partial match (PM) between LLMs’ generated answers and gold answers The prompts are shown in appendix A. The PM scores in table 3 and table 4 also reveal that LLMs are better at answering factual MHQA questions than counterfactual questions. In light of the results, we can find that LLMs achieve an inflated high performance on the Wikipedia-based MHQA dataset possibly because of the data contamination that leads to utilizing LLMs’ memory ability rather than reasoning ability.
Impact of Annotation on LLMs
@@
Model HotpotQA 1st Stage 2nd Stage EM F1 EM F1 EM F1 O1-preview 78.35 86.91 71.55 78.36 68.18 74.69 GPT-4 71.35 78.62 65.62 70.47 63.49 67.32 GPT-3.5 64.58 72.36 60.74 66.82 57.94 60.16 text 56.67 61.45 52.43 55.87 49.65 52.88 GEMINI-Pro 63.28 68.36 57.94 60.33 55.46 58.39 Bing-Chat 61.72 65.16 53.78 57.69 51.97 54.37
Table 5: Impact on LLMs’ performance at different stages in the annotation framework.
Our annotation framework involves two stages: keywords replacement followed by paraphrasing. To demonstrate the impact of annotated passages in different stages, three experiments were conducted using the original passages, passages generated from the first stage, and passages generated from the second stage. We additionally select and annotate another 300 passages from the HotpotQA dataset. Additionally, to ensure consistency in question complexity, we only annotated 2-hop questions on both sets of passages, aligning with the HotpotQA dataset.
Results are illustrated in Table 5, we find that: 1) there is also an obvious performance gap of LLMs between the original HotpotQA data and our two-stage data. Taking GPT-4 as an example, GPT-4 performs better on HotpotQA than our two-stage data, EM score of 71.35 and F1 score of 78.62 respectively. In the 1st stage data, GPT-4 shows a significant decrease, 65.62 EM and 70.47 F1 respectively. 2) The two stages of bias bring the different extents of performance drop. Comparing the performance of GPT-4 in stages 1 and 2, there is also an obvious drop, from 65.62 to 63.49 and from 70.47 to 67.32. It is mainly because of the data produced from the first stage although some keywords such as named entities and noun phrases, other words, sentence structure, and internal logic are the same as the original passages. The second stage passages are totally different from the original passages. Table 10 also illustrates an example from HotpotQA that is annotated in different stages. We find that the passage in stage 1 still maintains the same sentence structure. While the passage in stage 2 is totally different from the original passage.
@@
Figure 3 illustrates the performance degradation because LLMs also suffer from error propagation with the reasoning depth increasing. Figure 4 shows the performance of LLMs on the different hops of questions. According to the observation of the three figures, we find that with the hop increases, the complexity of multi-hop questions also increases, leading to the LLMs’ performance decrease. When incorrectly answering the previous sub-question, the latter one will also be influenced. Consequently, the performance of Sub_Question 2 is worse than that of Sub_Question 1. Tables 13 and 14 also illustrate the sub-qa performance of LLMs on the 3-hop and 4-hop datasets in appendix C.
Setting 2 hop 3 hop 4 hop EM F1 EM F1 EM F1 GPT-4 w/o Sub-Q 43.8±0.2 65.2±0.3 41.4±0.2 61.6±0.4 38.1±0.5 48.9±0.3 w Sub-Q 53.2±0.5 67.7±0.6 44.5±0.2 64.5±0.3 42.1±0.1 53.1±0.4 GPT-3.5 w/o Sub-Q 34.3±0.2 51.3±0.1 32.7±0.3 48.6±0.3 31.2±0.5 41.7±0.4 w sub-Q 40.4±0.5 56.9±0.4 37.5±0.1 50.2±0.2 33.5±0.2 45.9±0.6 GEMINI-pro w/o Sub-q 25.2±0.5 55.2±0.7 20.8±0.6 38.7±0.4 14.1 ±0.2 31.0±0.4 w sub-Q 34.6±0.5 64.2±0.5 27.3±0.2 42.1±0.2 25.9±0.2 33.8 ±0.3 text-davinci-003 w/o Sub-q 24.1±0.7 48.8±0.3 22.1±0.4 45.6±0.2 20.0 ±0.1 42.1±0.2 w sub-Q 32.3±0.4 52.7±0.4 27.3±0.3 46.4±0.3 24.2±0.5 42.8±0.7 Bing Chat w/o Sub-q 37.2±0.2 52.4±0.5 33.3±0.5 44.2±0.5 29.3±0.4 38.7±0.3 w sub-Q 41.8±0.5 56.8±0.5 40.1±0.2 48.4±0.5 32.4±0.3 41.3±0.3 O1-preview w/o Sub-q 52.3±0.3 65.4±0.4 49.3±0.5 60.2±0.5 42.5±0.3 55.9±0.2 w sub-Q 56.7±0.3 72.3±0.5 40.1±0.2 65.4±0.5 46.4±0.3 60.7±0.3
Table 6: Ablation study of the MHQA task, where we remove the sub-question information from the prompt and only ask LLMs to give the final answer. Here all the prompts are in a Chain-of-Thought setting.
@@
We calculate the proportion of the reasoning chain on the 2-hop dataset and follow the settings of (Tang et al., 2021) in calculating the percentage of correct or incorrect answers and record the results.
Table 8 shows the reasoning chain evaluation results. The green row shows the percentage of examples whose multi-hop questions can be correctly answered from the right reasoning chain. The red rows show the percentage of examples whose multi-hop questions can be correctly answered but through an incorrect reasoning chain. Among these examples, we notice that there is a low percentage of the correct final answer based on the right reasoning chain. There is also a large proportion of incorrect final answers as shown in rows 2,4,6 and 8.
Taking the results of GPT-3.5 as an example, the right reasoning chain only accounts for 13.3% although it shows a relatively high QA performance in previous tables. The percentage of incorrect reasoning chain of GPT-3.5 is 17.7% (sum of the three red rows). However, total failure cases account for 69% (sum of rows 2, 4, 6, and 8) which is substantial for the whole dataset. We conclude that LLMs only get a small proportion of the right reasoning chain and their high performance is relatively inflated due to the considerable proportion of incorrect reasoning chain.
2 hop 3 hop 4 hop F1 $RC$ EM $RC$ F1 $RC$ EM $RC$ F1 $RC$ EM $RC$ O1-preview 0.6 1.3 0.9 1.3 1.9 3.8 GPT-4 0.7 1.5 1.1 1.6 2.3 4.2 GPT 3.5 1.7 2.4 2.7 3.2 3.6 5.8 GEMINI-pro 2.1 3.9 4.6 8.7 5.4 9.5 text-davinci-003 2.4 2.9 3.9 5.2 5.5 7.4 Bing Chat 0.9 1.9 4.2 8.4 4.7 8.9
Table 7: LLMs’ joint performance on the whole reasoning chain. The scores are the average scores of three experiment results. The larger score means a worse performance on the whole reasoning chain.
@@
The joint F1 $RC$ and joint EM $RC$ scores in Table 7 are the whole reasoning chain evaluation results. We find that with the increases in the reasoning chain, the performances of LLMs dropped swiftly. For example, Bing Chat gets comparable performance with GPT-4 (0.7 joint F1) on answering 2 hop questions and gets a 0.9 joint F1 score. However, in the 3-hop question, the joint F1 $RC$ and joint EM $RC$ scores of Bing Chat are 4.2 and 8.4. In the 4-hop dataset, Bing Chat gives 4.7 joint F1 $RC$, and 8.9 joint EM $RC$ scores, respectively. Since the joint performance is a negative log, the larger scores mean the worse performance on the reasoning chain. We can conclude that LLMs’ reasoning ability decreases with the increases in reasoning chain length.
$q_{sub1}$ $q_{sub2}$ $q$ O1-preview GPT-4 GPT-3.5 GEMINI-pro text Bing Chat c c c 45.2 36.3 13.3 15.0 17.3 28.3 c c w 9.8 12.3 9.3 9.0 10.7 7.7 c w c 3.2 2.0 6.7 5.3 7.7 6.0 c w w 15.4 25.3 24.3 14.7 25.0 16.3 w c c 4.9 5.7 3.7 5.3 6.7 2.3 w c w 4.1 3.7 3.7 5.3 3.7 3.0 w w c 0.8 0.3 7.3 13.3 8.7 5. 0 w w w 16.6 14.3 31.7 32.3 30.3 31.3
@@
Table 8: Categorical EM statistics (%) of sub-question evaluation for the five LLMs on our 2-hop dataset.c stands for correct and w stands for wrong. For example, the third row shows the percentage of questions where models correctly answer both 2-hop questions and the first sub-question but incorrectly answer the second sub-question. We abbreviate text-davinci-003 as text.
4.3 Impact of Sub-Question